OpenStudy (anonymous):

question related to Regression analysis.. find byx and bxy from the regression equations 4y=9x + 15 and 25x=6y + 7. after arranging and converting the equations to the standard form Y= a + byx.X and X = a' + bxy.Y , how do i know which equation is which? i.e. which equation is for byx and bxy?

OpenStudy (ranga):

Since the first equation is given as 4y=9x + 15 you can make the assumption this is the y on x equation and the 25x=6y + 7 as the x on y and state the assumptions in your answer.

OpenStudy (anonymous):

how can i be sure if my assumption is correct? what do i do if my assumed byx.bxy is more than 1?

OpenStudy (ranga):

Perhaps that will be a clue to reverse the assumption?

OpenStudy (anonymous):

ok got it! thanks byx=9/4 and bxy=6/25. which gives me Correlation(x,y)= bxy.byx = 0.54

OpenStudy (ranga):

yw.

OpenStudy (ranga):

Isn't the coefficient of correlation square root of bxy.byx? r = sqrt(0.54) = 0.73 ?